the length of a rectangle is 4meters greater than its width. the area is 96 meters square. what is the length and width
Answer by checkley77(12844)

L=W+4
A=LW
96=(W+4)W
96=W^2+4W
W^2+4W-96=0
(W-8)(W+12)=0
W-8=0
W=8 ANS. FOR THE WIDTH.
L=8+4=12 ANS. FOR THE LENGTH.
PROOF:
96=8*12
96=96